1. use the bullet information below to find the

a. Cost of debt

b. Cost of equity (use CAPM)

c. The current stock price (use DGM)

d. Total market value of equity

e. Total market value of debt

f. WACC

*the firm has 30 million shares of common stock outstanding.The beta is 2.

* The firm most recently paid $5.50 dividend and plans to increase dividends by 2% per year indefinitely.

* The expected return on the market portfolio is 12% and the risk-free rate is 4%

*The firm has one bond issue outstanding. All bonds make annual coupon payments and are 5% coupon bonds with the face value of $1000. They are currently trading at face value. There are 100,000 of these bonds.

*The firms average tax rate is 30%

thank you!
Answer #1

Answers

1.Cost of debt

=Interest (1-Tax Rate)

= 5(1-0.30)

= 3.50%

Hence total cost of debt= $1000*100,000= $35 Million

2. Cost of equity using CAPM Approch

Ke= Rf+Beta(Rm-Rf)

= 4+2(12-4)

= 4+16

Ke = 20%

Where,

Ke= Cost of equity,

Rf= Risk free rate,

Rm= Return on market portfolio.

3.Current stock price using DGM Approach

P= D(1+g)/Ke-g

P= $5.50(1.02)/0.20-0.02

P= $5.61/0.18

P= $31.17

Where,

D= Dividend, P= Price of stock, g=Growth Rate

4.Market value of equity

=Current market price per share* No. of shares

=$31.17*30 million

=$935.10 million

5.Maket value of debt

= $1000*100,000= $100 million

6. WACC

Sr. No. Sources Value(million) Weight Cost Weight* Cost
a Equity $935.10 0.9034 20% 18.07%
b Debt $100 0.966 3.5% 0.34%
Total 18.41%

Note- WACC has been calculated using market value approch, we can also calculate using book value approch.